OpenStudy (anonymous):

HELP! Write an equation in standard form that passes through points (-4,10) and (0,-5).

OpenStudy (anonymous):

A. 15x-4y=20 B. 4x-15y=20 C. 15x+4y= -20 D. 4x+15y= -20

OpenStudy (dumbcow):

can you find the slope of the line?

OpenStudy (dumbcow):

once you have the slope, equation of line is: \[y - y_1 = m(x - x_1)\] \[(x_1 , y_1) = (0,-5)\]

OpenStudy (anonymous):

I'm confused.

OpenStudy (anonymous):

How do I find the slope?

OpenStudy (theopenstudyowl):

Hello @cheymariee , do you still need some help?

OpenStudy (anonymous):

Yes I do.

OpenStudy (en):

you will find the slope using \[m=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] just substitute you points.. so lets say that your first x coordinate is 0, and your first y coordinate is -5. the points (-4,-10) will be your second x and y coordinate...

OpenStudy (dumbcow):

\[slope = m = \frac{y_2 - y_1}{x_2 - x_1}\] rise over run

OpenStudy (anonymous):

-5/-4